### v2

@@ 3.2 Increasing Grace Period
The increasing grace period $[t_{1}(i),t_{2}(i)]$ for type $i$ customers is defined as: for any customer $i(k)$ that arrives within the interval from $t_{1}(i)$ to $t_{2}(i)$, customer $i(k)$ is accepted with probability $\alpha$ if customer $i(k-1)$ was rejected, and customer $i(k)$ is accepted if customer $i(k-1)$ was accepted.
@@ 4 Individually Fair RM Algorithms under Stochastic Arrivals
In this section, we focus on making three algorithm structures mentioned in Section 2.2DLP-PA, RDLP-PA, and S-BPC– individually fair without incurring significant additional regret. Based on the techniques we use for these three algorithms, we also enhance a more advanced algorithm – Bid Price Control with Online Gradient Descent (BPC-OGD) to satisfy the fairness metrics in Appendix 13. Further, as Algorithm DLP-PA is a variant for RDLP-PA (resolve 0 times), we analyze DLP-PA and RDLP-PA together.
@@ 4.1 Algorithm DLP-PA and RDLP-PA
Here, we introduce the Grace Period Enhanced RDLP-PA algorithm (GP-Enhanced-RDLP). The regret of the original RDLP-PA algorithm is $O(T^{\beta})$, where $\beta$ is a parameter that depends on the frequency of resolving the DLP. For example, $\beta=1/2$ corresponds to resolving zero times (i.e., Algorithm DLP-PA), while $\beta=1/3$ corresponds to resolving one or more times.
We begin by uniformly partitioning the time horizon $[0,T]$ into $K=T^{1-\beta}$ segments of size $T^{\beta}$. We solve the DLP with the initial capacity vector at $t=0$, and solve the DLP with the remaining capacity vector at $t=t^{*}$, where $t^{*}$ is the re-solving time point of the original RDLP-PA algorithm. In the first time segment ($k=1$), for each type $i$ customer, we sample a random variable $y_{i}\sim\text{Bin}(\lambda_{i}T^{\beta},\frac{x_{i}^{*}}{\lambda_{i}})$. Our objective is to accept approximately $y_{i}$ customers of type $i$ in this interval, akin to the original RDLP-PA algorithm’s acceptance of $\text{Bin}(\lambda_{i}T^{\beta},\frac{x_{i}^{*}}{\lambda_{i}})$ customers of type $i$. However, different from RDLP-PA’s probabilistic allocation, we initially accept $y_{i}-\bar{a}n\gamma-1$ customers of type $i$, where $\gamma=\log_{1-\alpha}\delta$. Subsequently, we implement a decreasing grace period for type $i$ customers until the end of the first interval. According to Theorem 3.2, this ensures the individual fairness metrics while accepting $z_{i}^{(1)}=\Theta(\log T)$ fewer customers of type $i$ compared to Algorithm RDLP-PA.
Starting from the second time segment $(k\geq 2)$, due to the probable rejection of the last few customers in segment $k-1$, an immediate acceptance of customers in segment $k$ is unfair. Thus, we introduce an increasing grace period, progressively shifting from rejection to acceptance. Notably, in segment $k-1$, as $z_{i}^{(k-1)}$ fewer customers of type $i$ are accepted compared to Algorithm RDLP-PA, we compensate by accepting $z_{i}^{(k-1)}$ additional type $i$ customers in segment $k$. This is achieved by drawing $y_{i}\sim z_{i}^{(k-1)}+\text{Bin}(\lambda_{i}T^{\beta},\frac{x_{i}^{*}}{\lambda_{i}})$ and implementing an increasing grace period until the arrival of the $(y_{i}-\bar{a}n\gamma)^{\text{th}}$ customer of type $i$. Subsequently, a decreasing grace period is applied until the conclusion of the time segment.
@@
Input: Preference matrix A. Arriving rate vector $\mathbf{\lambda}$. Capacity vector m. Resolving time point $t^{*}$. Regret bound of input R-DLP algorithm $O(T^{\beta})$. Fairness parameters $(\alpha,\delta)$. $\gamma=\log_{1-\alpha}\delta$.
@@
Initialize: Vector $\mathbf{z}^{(0)}=0$. $K=T^{1-\beta}$.
@@
for $k\in\{1,2,...,\lfloor t^{*}T^{-\beta}\rfloor\}$ do
       Run Algorithm 2 with $t_{1}=(k-1)T^{\beta}$, $t_{2}=kT^{\beta}$, $\mathbf{z}=\mathbf{z}^{(k-1)}$, $\mathbf{x}^{*}$. Get the output $\mathbf{z}^{(k)}\leftarrow\mathbf{w}$.
@@
for $k\in\{\lceil t^{*}T^{-\beta}\rceil,\lceil t^{*}T^{-\beta}\rceil+1,\ldots,K\}$ do
       Run Algorithm 2 with $t_{1}=(k-1)T^{\beta}$, $t_{2}=kT^{\beta}$, $\mathbf{z}=\mathbf{z}^{(k-1)}$, $\tilde{\mathbf{x}}^{*}$. Get the output $\mathbf{z}^{(k)}\leftarrow\mathbf{w}$.
@@
Given any $\alpha,\delta\in(0,1)$, and any input RDLP-PA algorithm that attains a regret of $O(T^{\beta})$ which resolves $K$ times of LP, Algorithm GP-Enhanced-RDLP: (i) is $(\alpha,\delta)$-fair, and (ii) incurs regret $O(\min\{T^{\beta},K\log 1/\delta\})$. Specifically, for $\delta\geq 1/\left(\text{exp}(T^{\beta}-e^{K}\right)$, the regret is $O(T^{\beta})$. [^4] For the state-of-the-art Algorithm (Bumpensanti and Wang 2020), $K=O(\log\log T)$, and the lower bound of $\delta$ is very close to $\exp(-T^{\beta})$. Recall that $1-\delta$ is the confidence level of the individual fairness metrics, here our result shows that no addition asymptotic regret is incurred if the confidence level is no more than $1-\exp(-T^{\beta})$. For large $T$, the confidence level is very close to $1$.
@@
Here, we provide a proof sketch for $K=1$, where Rev(RDLP-PA)re-solves the LP once. Algorithm GP-Enhanced-RDLP is $(\alpha,\delta)$-fair since it introduces a grace period every time that a decision transitions from acceptance to rejection and vice versa. To confirm that the regret of Algorithm GP-Enhanced-RDLP is $O(T^{\beta})$, we have
@@
$=O(T^{\beta})+\text{Rev}(\textsf{RDLP-PA})-\text{Rev}(\pi),$
@@
Let $\tau$ be a random variable indicating the stopping time at which some resource becomes exhausted (triggering the termination of the algorithm), with $\tau$ belonging to time segment $k(\tau)$. Ignoring the grace period and accepting $y_{i}$ type $i$ customers in each segment, the expected revenue of Algorithm GP-Enhanced-RDLP almost matches that of Algorithm RDLP-PA for $t\in[0,(k(\tau)-1)T^{\beta}]$.
Furthermore, although the application of both increasing and decreasing grace periods in each time segment may lead to the additional loss of up to $\Theta(\log 1/\delta)$ customers of each type, we accept an equivalent number of customers of each type in the following time segment. By doing this, before $t^{*}$, the expected difference in the number of accepted customers of each type between Algorithm RDLP-PA and Algorithm GP-Enhanced-RDLP stems only from the last time segment before $t^{*}$, amounting to $\Theta(\log 1/\delta)$. Therefore, the expected difference in remaining capacity between Algorithm RDLP-PA and Algorithm GP-Enhanced-RDLP is bounded by $\Theta(\log 1\delta)$ at time $t^{*}$. As such, the disparity in the re-solved optimal solutions at $t^{*}$ between Algorithm RDLP-PA and Algorithm GP-Enhanced-RDLP is negligible. Consequently, we eliminate all revenue loss due to the grace period for all $k(\tau)-1$ time segments. For the time segment $k(\tau)$ of size $T^{\beta}$, we may lose a maximum of $O(T^{\beta})$ in revenue compared to the RDLP-PA algorithm. Hence, we have
@@ 4.2 Algorithm S-BPC
Given any $\alpha,\delta\in(0,1)$, and any input S-BPC algorithm with a regret of $O(\sqrt{T})$, by giving a decreasing grace period $[\min_{j\in[L]}m_{j}-\bar{a}n\gamma,T]$ to each type $i$ customer (for $i$ such that $r_{i}>\sum_{j=1}^{L}\theta_{j}^{*}A_{ij}$): (i) the algorithm is $(\alpha,\delta)$-fair, (ii) incurs regret $O(\min\{\sqrt{T},\log 1/\delta\})$. Specifically, for $\delta\geq\text{exp}(-\sqrt{T})$, the regret is $O(\sqrt{T})$.

### v3

@@ 4.2 Increasing Grace Period
The increasing grace period $[t_{1}(i),t_{2}(i)]$ for type $i$ customers is defined as: for any customer $i(k)$ that arrives within the interval from $t_{1}(i)$ to $t_{2}(i)$, customer $i(k)$ is accepted with probability $\beta$ if customer $i(k-1)$ was rejected, and customer $i(k)$ is accepted if customer $i(k-1)$ was accepted.
@@ 5 Individually Fair RM Algorithms under Stochastic Arrivals
In this section, we focus on making four algorithm structures mentioned in Section 2.2 (DLP-PA, RDLP-PA, S-BPC, BPC-OGD) individually fair without incurring significant additional regret. Further, as Algorithm DLP-PA is a variant for RDLP-PA (resolve 0 times), we analyze DLP-PA and RDLP-PA together. All primal and dual methods can be adapted to our revised algorithm structure since we focus only on the (re)solving of the primal/dual LP, rather than the specific decision-making process used by the original algorithms. That is, as long as different algorithms resolve the primal/dual LP at the same time, our revised approach remains consistent, regardless of their individual decision structures.
@@ 5.1 Algorithm DLP-PA and RDLP-PA
Here, we introduce the Grace Period Enhanced RDLP-PA algorithm (GP-Enhanced-RDLP). The regret of the original RDLP-PA algorithm is $O(T^{\eta})$, where $\eta$ is a parameter that depends on the frequency of resolving the DLP. For example, $\eta=1/2$ corresponds to resolving zero times (i.e., Algorithm DLP-PA), while $\eta=1/3$ corresponds to resolving one or more times.
We begin by uniformly partitioning the time horizon $[0,T]$ into $K$ segments. Each time segment has length $\max\{T^{\eta},2\bar{a}n\gamma\}$, where $\gamma=\log_{1-\beta}\delta$. $\bar{a}n\gamma$ is the length of the grace period. Then, we have $K=\min\{T^{1-\eta},T/(\bar{a}n\gamma)\}$. We split the discussion into two cases: (1) $T^{\eta}>2\bar{a}n\gamma$; (2) $T^{\eta}\leq 2\bar{a}n\gamma$.
Case 1: $T^{\eta}>\bar{a}n\gamma$. In this case, the regret is greater than the grace period length, which means that we are not resolving for many times. We illustrate this case by discussing the resolving once algorithm. We solve the DLP with the initial capacity vector at $t=0$, and solve the DLP with the remaining capacity vector at $t=t^{*}$, where $t^{*}$ is the re-solving time point of the original RDLP-PA algorithm. In the first time segment ($k=1$), for each type $i$ customer, we sample a random variable $y_{i}\sim\text{Bin}(\lambda_{i}T^{\eta},\frac{x_{i}^{*}}{\lambda_{i}})$. Our objective is to accept approximately $y_{i}$ customers of type $i$ in this interval, akin to the original RDLP-PA algorithm’s acceptance of $\text{Bin}(\lambda_{i}T^{\eta},\frac{x_{i}^{*}}{\lambda_{i}})$ customers of type $i$. However, different from RDLP-PA’s probabilistic allocation, we initially accept $y_{i}-\bar{a}n\gamma-1$ customers of type $i$, where $\gamma=\log_{1-\beta}\delta$. Subsequently, we implement a decreasing grace period for type $i$ customers until the end of the first interval. According to Theorem 4.2, this ensures the individual fairness metrics while accepting $z_{i}^{(1)}=\Theta(\log T)$ fewer customers of type $i$ compared to Algorithm RDLP-PA.
Starting from the second time segment $(k\geq 2)$, due to the probable rejection of the last few customers in segment $k-1$, an immediate acceptance of customers in segment $k$ is unfair. Thus, we introduce an increasing grace period, progressively shifting from rejection to acceptance. Notably, in segment $k-1$, as $z_{i}^{(k-1)}$ fewer customers of type $i$ are accepted compared to Algorithm RDLP-PA, we compensate by accepting $z_{i}^{(k-1)}$ additional type $i$ customers in segment $k$. Here, $z_{i}^{(k)}$ are independently distributed for different $k$, but in segment $k$, we can use the realization of $z_{i}^{(k-1)}$ to draw $y_{i}\sim z_{i}^{(k-1)}+\text{Bin}(\lambda_{i}T^{\eta},\frac{x_{i}^{*}}{\lambda_{i}})$ and implementing an increasing grace period until the arrival of the $(y_{i}-\bar{a}n\gamma)^{\text{th}}$ customer of type $i$. Subsequently, a decreasing grace period is applied until the conclusion of the time segment.
@@
Case 2: $T^{\eta}\leq\bar{a}n\gamma$. Suppose the resolving algorithm achieves a regret smaller than the grace period length, such as the $O(1)$ regret bound shown in Vera and Banerjee (2019), Bumpensanti and Wang (2020). In this case, we partition the time horizon into segments of length at least twice the grace period length. However, if the algorithm resolves too frequently—e.g., Vera and Banerjee (2019) resolves the DLP at every time step—this frequency is too high for a grace period to be maintained.
To address this, we accept a modest sacrifice in revenue in order to maintain the grace period design and enforce the fairness constraint. Consider, for example, the algorithm proposed by Vera and Banerjee (2019), which resolves the deterministic linear program (DLP) at every time step. Our enhanced version only resolves the DLP at periodic time points $\eta\bar{a}n\gamma$, where $\eta\in\left[T/(\bar{a}n\gamma)\right]$. This modification reduces the computational complexity and ensures fairness through the grace period design, albeit with the possibility of slightly reduced performance compared to the original algorithm. Nevertheless, similar to the analysis in Case 1, the regret of the enhanced algorithm remains bounded by the grace period length $O(\bar{a}n\gamma)=O(\log_{1-\beta}\delta)$.
Input: Preference matrix A. Arriving rate vector $\mathbf{\lambda}$. Capacity vector m. Resolving time point $t^{*}$. Regret bound of input R-DLP algorithm $O(T^{\eta})$. Fairness parameters $(\alpha,\delta)$. $\gamma=\log_{1-\beta}\delta$.
@@
Initialize: Vector $\mathbf{z}^{(0)}=0$. $K=T/(\max\{T^{\eta},2\log_{1-\beta}\delta\})$.
@@
for $k\in\{1,2,...,\lfloor t^{*}/(\max\{T^{\eta},2\log_{1-\beta}\delta\})\rfloor\}$ do
       Run Algorithm 2 with $t_{1}=(k-1)(\max\{T^{\eta},2\log_{1-\beta}\delta\})$, $t_{2}=k(\max\{T^{\eta},2\log_{1-\beta}\delta\})$, $\mathbf{z}=\mathbf{z}^{(k-1)}$, $\mathbf{x}^{*}$. Get the output $\mathbf{z}^{(k)}\leftarrow\mathbf{w}$.
@@
for $k\in\{\lceil t^{*}/(\max\{T^{\eta},2\log_{1-\beta}\delta\})\rceil,\lceil t^{*}/(\max\{T^{\eta},2\log_{1-\beta}\delta\})\rceil+1,\ldots,K\}$ do
       Run Algorithm 2 with $t_{1}=(k-1)(\max\{T^{\eta},2\log_{1-\beta}\delta\})$, $t_{2}=k(\max\{T^{\eta},2\log_{1-\beta}\delta\})$, $\mathbf{z}=\mathbf{z}^{(k-1)}$, $\tilde{\mathbf{x}}^{*}$. Get the output $\mathbf{z}^{(k)}\leftarrow\mathbf{w}$.
@@
Given any $\alpha,\delta\in(0,1)$, and any input RDLP-PA algorithm that attains a regret of $O(T^{\eta})$, Algorithm GP-Enhanced-RDLP: (i) is $(\alpha,\delta)$-fair, and (ii) incurs regret $O(\max\{T^{\eta},\log_{1-\beta}\delta\})$.
Remark 5.2
Specifically, when setting $\delta=1/T$, if RDLP-PA incurs regret $T^{\eta}=\omega(\log T)$, then GP-Enhanced-RDLP maintains the same regret bound of $T^{\eta}$. Conversely, if RDLP-PA has regret $T^{\eta}=O(\log T)$, then GP-Enhanced-RDLP achieves a regret bound of $\Theta(\log T)$. This result shows that the revenue loss introduced by GP-Enhanced-RDLP remains minimal.
@@
Here, we provide a proof sketch for the case where Rev(RDLP-PA) re-solves the LP once. Algorithm GP-Enhanced-RDLP is $(\alpha,\delta)$-fair since it introduces a grace period every time that a decision transitions from acceptance to rejection and vice versa. To confirm that the regret of Algorithm GP-Enhanced-RDLP is $O(T^{\eta})$, we have
@@
$=O(T^{\eta})+\text{Rev}(\textsf{RDLP-PA})-\text{Rev}(\pi),$
@@
Let $\tau$ be a random variable indicating the stopping time at which some resource becomes exhausted (triggering the termination of the algorithm), with $\tau$ belonging to time segment $k(\tau)$. Ignoring the grace period and accepting $y_{i}$ type $i$ customers in each segment, the expected revenue of Algorithm GP-Enhanced-RDLP almost matches that of Algorithm RDLP-PA for $t\in[0,(k(\tau)-1)T^{\eta}]$.
Furthermore, although the application of both increasing and decreasing grace periods in each time segment may lead to the additional loss of up to $\Theta(\log 1/\delta)$ customers of each type, we accept an equivalent number of customers of each type in the following time segment. By doing this, before $t^{*}$, the expected difference in the number of accepted customers of each type between Algorithm RDLP-PA and Algorithm GP-Enhanced-RDLP stems only from the last time segment before $t^{*}$, amounting to $\Theta(\log 1/\delta)$. Therefore, the expected difference in remaining capacity between Algorithm RDLP-PA and Algorithm GP-Enhanced-RDLP is bounded by $\Theta(\log 1\delta)$ at time $t^{*}$. As such, the disparity in the re-solved optimal solutions at $t^{*}$ between Algorithm RDLP-PA and Algorithm GP-Enhanced-RDLP is negligible. Consequently, we eliminate all revenue loss due to the grace period for all $k(\tau)-1$ time segments. For the time segment $k(\tau)$ of size $T^{\eta}$, we may lose a maximum of $O(T^{\eta})$ in revenue compared to the RDLP-PA algorithm. Hence, we have
@@ 5.2 Algorithm S-BPC
Given any $\alpha,\delta\in(0,1)$, and any input S-BPC algorithm with a regret of $O(\sqrt{T})$, by giving a decreasing grace period $[\min_{j\in[L]}m_{j}-\bar{a}n\gamma,T]$ to each type $i$ customer (for $i$ such that $r_{i}>\sum_{j=1}^{L}\theta_{j}^{*}A_{ij}$): (i) the algorithm is $(\alpha,\delta)$-fair, (ii) incurs regret $O(\max\{\sqrt{T},\log_{1-\beta}\delta\})$.
Remark 5.4
Specifically, for $\delta=O\left((1-\alpha)^{\sqrt{T}}\right)$, the regret is $O(\sqrt{T})$.
5.3 Algorithm BPC-OGD
In the Algorithm BPC-OGD, instead of using a fixed bid price, it dynamically updates the bid price with some first order optimization methods, such as online gradient descent or online mirror descent. The details of BPC-OGD can be found in Algorithm 9. BPC-OGD can be markedly unfair. To see this, let us simplify the setting and assume that there is only one type of customer for $t=1,2,\ldots,T$. Also assume that there is only one type of resource and each arriving customer requests a single unit of this resource. Subsequently, if we follow steps 15 and 16 in Algorithm 9, we find that $\theta^{(t)}$ is a deterministic process that can be described using the following formula:
$\theta^{(t+1)}$ =$=\left\{\begin{matrix}\;\theta^{(t)}-\frac{D}{G\sqrt{T}}\frac{m(t)}{T}+\frac{D}{G\sqrt{T}},&\text{ if $\theta^{(t)}<r$ },&\\ \;\theta^{(t)}-\frac{D}{G\sqrt{T}}\frac{m(t)}{T},&\text{ if $\theta^{(t)}\geq r$ },&\\ \end{matrix}\right.$ (8)
Here, $r\in(0,1)$, $m(t)$ denotes the number of resources remaining at time $t$, with $m(0)=m$, and
$m(t+1)$ =$=\left\{\begin{matrix}\;m(t)-1,&\text{ if $\theta^{(t)}<r$ },&\\ \;m(t),&\text{ if $\theta^{(t)}\geq r$ },&\\ \end{matrix}\right.$
The deterministic process $\theta^{(t)}$, as defined in Equation (8), allows for the precise computation of how frequently decisions switch between accepting and rejecting adjacent customers. Regardless of the constants $D$, $G$, and $r$, this decision-making flips $\Theta(T)$ times between acceptance and rejection. This behavior highlights the algorithm’s significant unfairness, as it implies that almost every customer is treated differently compared to their immediate predecessor or successor.
To modify BPC-OGD to comply with $(\alpha,\delta)$-fairness while minimally impacting revenue, we first reject all arrivals in the period $t\in[0,T^{2/3}]$. Meanwhile, we operate BPC-OGD as an auxiliary algorithm, documenting its decisions. We designate $u_{i}$ as the number of accepted type $i$ customers by BPC-OGD within the $[0,T^{2/3}]$ time horizon, and $\Lambda_{i}$ as the count of arriving type $i$ customers within $[0,T^{2/3}]$. Then, we obtain that the Euclidean distance between the rate of accepted type $i$ customers by BPC-OGD ($\frac{u_{i}}{\Lambda_{i}}$), and that by DLP-PA ($\frac{x_{i}^{*}}{\lambda_{i}}$), is bounded with high probability. Thus, even without access to the $\lambda_{i}$ arrival rate and with the prohibition of linear programming, we can utilize the initial $T^{2/3}$ periods to approximate the optimal DLP solution $\textbf{x}^{*}$. Finally, for $t\in[T^{2/3},T]$, we execute the GP-Enhanced-RDLP Algorithm with a $T^{2/3}$ time segment length, setting $x_{i}^{*}=\frac{u_{i}}{T^{2/3}}$ and $\lambda_{i}=\frac{\Lambda_{i}}{T^{2/3}}$, culminating in a $(\alpha,\delta)$-fair algorithm with regret $O(T^{2/3}\log T)$ if $T^{2/3}\log T=\Omega(\log_{1/(1-\beta)}1/\delta)$. Otherwise, the regret is the length of the grace period, which is $O(\log_{1-\beta}\delta)$. Full details can be found in Algorithm 3, and the following theorem summarizes the results. The proof can be found in Appendix 15.
Input: Preference matrix A. Reward vector r. Capacity vector m. OGD Parameter: $G$, $D$, $\bar{\theta}$. Fairness parameters $(\alpha,\delta)$.
Output: Update of dual variable $\mathbf{\theta}$ and acceptance decisions over time.
Initialize: Dual variable $\mathbf{\theta}^{(0)}\leftarrow\mathbf{0}$.
Reject all customers arriving between $t\in[0,\max\{T^{2/3},2\log_{1-\beta}\delta\}]$.
Denote $\Lambda_{i}$ as the number of arriving customers of type $i$ between $t\in[0,\max\{T^{2/3},2\log_{1-\beta}\delta\}]$, $i\in[n]$.
Run Algorithm BPC-OGD as an auxiliary algorithm for $t\in[0,\max\{T^{2/3},2\log_{1-\beta}\delta\}]$.
Denote $u_{i}$ as the number of accepted customers of type $i$ by BPC-OGD, $i\in[n]$.
Set $K=(T/\max\{T^{2/3},2\log_{1-\beta}\delta\})-1$, $\lambda_{i}=\frac{\Lambda_{i}}{\max\{T^{2/3},2\log_{1-\beta}\delta\}}$ and $x_{i}^{*}=\frac{u_{i}}{\max\{T^{2/3},2\log_{1-\beta}\delta\}}$, $\textbf{z}^{(0)}=\textbf{0}$.
for $k\in\{1,2,...,K\}$ do
       Run Algorithm 2 with $t_{1}=k\max\{T^{2/3},2\log_{1-\beta}\delta\}$, $t_{2}=(k+1)\max\{T^{2/3},2\log_{1-\beta}\delta\}$, $\mathbf{z}=\mathbf{z}^{(k-1)}$, optimal primal solution $\mathbf{x}^{*}$. Denote the output vector $\mathbf{w}$ as $\mathbf{z}^{(k)}$.
Algorithm 3 GP-Enhanced-BPC-OGD
Theorem 5.5
Given any $\alpha,\delta\in(0,1)$, and any input BPC-OGD algorithm that attains a regret of $O(\sqrt{T})$, under a mild non-degenerate assumption (assumption 3 in Agrawal et al. (2014)), Algorithm GP-Enhanced-BPC-OGD: (i) is $(\alpha,\delta)$-fair, (ii) incurs regret $O(\max\{T^{2/3}\log T,\log_{1-\beta}\delta\})$.
Remark 5.6
For $\delta=O\left((1-\beta)^{T^{2/3}\log T}\right)$, the regret of GP-Enhanced-BPC-OGD is $O(T^{2/3}\log T)$.